OpenStudy (anonymous):

convert to a fraction in lowest terms. 16.4%

OpenStudy (anonymous):

41/250

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

16.4% is equal to 0.164, so that's the decimal you want your fraction to equal. (16.4/100)=(164/1000) then it's just a matter of simplfying. 164/4=41 and 1000/4 is equal to 250. since nominator and denominator cannot be reduced any further (into whole numbers), your answer is 41/250.

OpenStudy (anonymous):

thank u